# BianCang: A Traditional Chinese Medicine Large Language Model

### v1

@@ 3.2 Supervised Fine-Tuning (SFT)
TCM-Plan-SFT Medical Record Instruction 21M 1,3409 70MB
@@
SFT is the core stage for applying the large language model of TCM to real-world scenarios, and the key phase in enabling the model’s conversational abilities. The quality and diversity of the instruction data are crucial to the SFT process [Touvron et al., 2023]. To maximize the activation of the knowledge embedded in the pre-trained model and align it with downstream real-world TCM tasks, we constructed a high-quality and diverse instruction dataset, which contains the following five types of data:
@@
TCM Pharmacopoeia Data: Based on the Pharmacopoeia of the People’s Republic of China (Part I: Traditional Chinese Medicine), we constructed the ChP-KnowledgeQA and ChP-PrescriptionWriting instruction sets to enhance the model’s TCM knowledge and question-answering abilities. The ChP-KnowledgeQA set consists of question-answer pairs covering topics such as the properties, identification, processing, taste, meridian tropism, functions, indications, dosage, and storage methods of TCM herbs and medicinal materials. The ChP-PrescriptionWriting set consists of question-answer pairs focused on composing and formulating TCM prescriptions. When constructing these two instruction sets, we manually created several query templates to ensure the diversity of the instructions. These two datasets are referred to as the ChP-TCM dataset.
@@
Medical Record Data: Based on real hospital records, we constructed the TCMSD-DD-SFT, TCM-WM-Diff-SFT, and TCM-Plan-SFT instruction sets to improve the model’s performance on real-world tasks such as syndrome differentiation and disease diagnosis in TCM. The TCMSD-DD-SFT set is designed for TCM syndrome and disease differentiation, aiming to analyze a patient’s condition and syndrome type based on their chief complaint, admission details, medical history, and information gathered through TCM’s “Four Diagnoses” (inspection, listening and smelling, inquiry, and palpation). The TCM-WM-DiffDiag-SFT set includes differential diagnoses in both TCM and western medicine (WM) based on real hospital records. The TCM-Plan-SFT set focuses on providing detailed treatment plans based on a patient’s basic condition and diagnosis results. To construct these instruction sets, we manually created a variety of inquiry and response templates, and then employed a weighted random selection approach to generate instruction data from these templates, enhancing the quality and diversity of the instruction sets.
@@
Multi-turn Dialogue Data: To equip the model with conversational capabilities in the medical field, we utilized the DISC-Med-SFT dataset [Bao et al., 2023]. DISC-Med-SFT is a high-quality SFT dataset containing a large number of samples that were reconstructed from existing medical datasets. These data help the model learn medical knowledge, align its behavior with human preferences, and adapt to the distribution of real-world online medical dialogues.
@@
Exam Data: The MLEC-QA dataset is a large-scale Chinese biomedical multiple-choice question-answering dataset [Li et al., 2021], sourced from the Chinese Medical Licensing Examination. It covers five subfields: TCM, integrative medicine (combining TCM and Western medicine), clinical medicine, public health, and basic clinical medicine. Based on this dataset, we constructed the MLEC-SFT instruction set to enhance the model’s comprehensive question-answering abilities in TCM and healthcare.
@@
General Domain Data: To prevent the loss of the model’s general capabilities, we mixed in some general domain data during the SFT process. MS-Agent-Bench [^1] https://www.modelscope.cn/datasets/iic/MSAgent-Bench is a general-purpose multi-domain SFT dataset, from which we randomly selected 65,000 samples. Self-Cognition is a custom instruction set we manually constructed to help the model develop self-awareness capabilities.
@@
3.3 Model Training
We train BianCang at two scales, with parameter sizes of 7 billion and 14 billion, based on Qwen2/2.5-7B and Qwen2.5-14B [Yang et al., 2024a], respectively. We employ full-parameter fine-tuning during both the continuous pre-training and supervised fine-tuning stages. We use Scalable lightWeight Infrastructure for Fine-Tuning (SWIFT) [Zhao et al., 2024] as the training framework.
To ensure training stability, we set the weight decay to 0.1 for 14B model and 0.01 for 7B model to prevent overfitting, the max gradient norm to 0.5 to prevent gradient explosion, and the warmup ratio to 0.05 for 14B model and 0.03 for 7B model to maintain smooth training. To balance training costs, we utilize parallel training, gradient checkpointing, fp16 precision, and gradient accumulation strategy, and limit the model’s maximum response length to 4096 tokens. We reserve 5% of the training set for validation and use token-level accuracy to validate the model, saving the best checkpoint as the final model. Experiments are conducted on 8 A100 40G GPUs, with both continuous pre-training and supervised fine-tuning performed over two epochs. The loss across all training stages successfully converges to a valid range.
4 Experiments and Evaluation
@@ 4.1 Baselines
GPT-4 [Achiam et al., 2023]: GPT-4 is a large-scale, multimodal model developed by OpenAI. It has demonstrated human-level performance across various professional and academic benchmarks. We conduct tests using the API service provided by OpenAI, invoking the GPT-4o model.
@@
BianQue2-6B [Chen et al., 2023c]: BianQue-6B is a ChatGLM-based [Du et al., 2022] LLM finetuned with the self-constructed health conversation dataset BianQueCorpus that is consist of multiple turns of questioning and health suggestions polished by ChatGPT. BianQue2-6B expands upon the previous version by incorporating data from pharmaceutical instruction guides, medical encyclopedia knowledge, and ChatGPT distillation instructions, enhancing the model’s capabilities in providing recommendations and knowledge retrieval.
@@
DoctorGLM-6B [Xiong et al., 2023]: A large-scale Chinese medical model based on ChatGLM-6B [Du et al., 2022], fine-tuning on a large amount of medical instruction dataset.
@@
BenTsao-Llama-7B [Wang et al., 2023]: A Chinese medical LLM based on Chinese-LLaMA-7B [Cui et al., 2023], and fine-tuned on an 8K medical dialogue dataset.
@@
ShenNong-TCM-LLM-7B [Wei et al., 2023]: A large traditional Chinese medicine model based on Chinese-LLaMA-7B [Cui et al., 2023], fine-tuned on the self-constructed ChatMed-TCM dataset.
@@
Taiyi-7B [Luo et al., 2024a]: A bilingual medical large language model designed for various biomedical tasks, trained on 140 Chinese and English biomedical datasets across 15 BioNLP tasks.
@@
WiNGPT2222https://github.com/winninghealth/WiNGPT2: A large language model developed by Weining Health for the medical domain, aimed at providing intelligent medical Q&A, diagnostic support, and medical knowledge services to improve the efficiency of diagnosis and the quality of healthcare services. WiNGPT2-7B-Base is based on Qwen-7B [Bai et al., 2023], while WiNGPT2-Llama3-8B-Base is based on LLaMA3-8B [Dubey et al., 2024].
@@
HuatuoGPT2 [Chen et al., 2023a]: HuatuoGPT2 employs an innovative domain adaptation method to significantly boost its medical knowledge and dialogue proficiency. We utilized two models, HuatuoGPT2-7B and HuatuoGPT2-13B, which are based on Baichuan2-7B-Base and Baichuan2-13B-Base [Yang et al., 2023], respectively.
@@
Sunsimiao-7B [Xin and Dong, 2023]: A Chinese medical large language model based on Qwen2-7B [Yang et al., 2024a], fine-tuned on high-quality medical data. It has achieved excellent results in Chinese medical exams.
@@
QiZhen-CaMA-13B [^3] https://github.com/CMKRG/QiZhenGPT: A Chinese medical large language model based on CaMA-13B [Zhang et al., 2023b], trained on the QiZhen medical knowledge base with a Chinese medical instruction set. It performs exceptionally well in Chinese medical scenarios.
@@
Zhongjing-13B [Yang et al., 2024b]: A Chinese medical large language model that has undergone end-to-end training, from pre-training to reinforcement learning. In certain dialogue scenarios, it approaches the professional level of expert doctors. Zhongjing-13B is based on Ziya-LLaMA-13B-v1 [Zhang et al., 2022], a general Chinese LLM with 13 billion parameters, trained based on LLaMA.
@@
ChiMedGPT-13B [Tian et al., 2024]: ChiMed-GPT is a Chinese medical large language model based on Ziya-LLaMA-13B-v2 [Zhang et al., 2022]. It has undergone comprehensive pre-training, SFT, and RLHF.
@@
DISC-MedLLM-13B [Bao et al., 2023]: A large language model specifically designed for medical and healthcare conversational scenarios. DISC-MedLLM-13B is based on Baichuan-13B-Base [^4] https://github.com/baichuan-inc/Baichuan-13B and trained on the DISC-Med-SFT multi-turn dialogue dataset.
@@
Lingdan-13B [Hua et al., 2024]: A traditional Chinese medicine (TCM) large language model based on Baichuan2-13B-Base [Yang et al., 2023], trained on a rich dataset including TCM ancient books, textbooks, and the Chinese Pharmacopoeia.
@@
GLM4 [GLM et al., 2024]: GLM-4-9B is the open-source version of the latest generation of pre-trained models in the GLM-4 series launched by Zhipu AI.
@@
Qwen2/2.5 [Yang et al., 2024a]: Qwen2 is a series of open-source multilingual pre-trained and instruction-tuned models based on the Transformer architecture, developed by the Qwen Team. These models have achieved outstanding results across various benchmark evaluations. Qwen2.5 represents a significant upgrade over Qwen2, optimized to meet specific requirements.
4.2 Evaluation
We conducted model evaluation from both objective and subjective perspectives. The objective evaluation includes three tasks: TCM syndrome differentiation, TCM disease diagnosis, and exams, covering ten datasets. The subjective evaluation task involves medical record analysis using BC-Analytical test set. The prompt templates used in the experiments can be found in the appendix.
4.2.1 Objective Evaluation
For the TCM syndrome differentiation task, we used the TCMSD and TCMSD-BC test sets. The TCMSD test set was adapted from the test set of the TCM-SD dataset [Mucheng et al., 2022], where we constructed queries and responses using fields from the TCM-SD test set. The TCMSD-BC dataset was built based on proprietary medical records. The TCM syndrome differentiation task is divided into two test modes: direct inference (DI) and chain of thought (CoT), with accuracy serving as the evaluation metric.
For the TCM disease diagnosis task, we used the TCMDD and TCMDD-BC test sets. Similarly, the TCMDD test set was adapted from the test set of the TCM-SD dataset [Mucheng et al., 2022], while the TCMDD-BC dataset was constructed based on proprietary medical records. Like the syndrome differentiation task, the disease diagnosis task is evaluated using two test modes: DI and CoT, with accuracy as the evaluation metric.
To assess the model’s mastery of TCM and medical knowledge, we applied the MLEC-QA [Li et al., 2021] and Chinese Medical Benchmark (CMB) [Wang et al., 2024] test sets. The MLEC-QA test set is divided into five subsets: traditional Chinese medicine, traditional Chinese medicine combined with western medicine, clinic, public health, and stomatology. The CMB dataset is a comprehensive medical benchmark that includes six categories: physician exam, nursing exam, pharmacist exam, medical technician exam, professional knowledge exam, and medical postgraduate entrance exam. The average score across these six categories is used as the final result. The examination tasks are evaluated in two test modes: zero-shot and few-shot, with accuracy as the evaluation metric.
@@ 4.2.1 Objective Evaluation
GPT-4 24.53 45.21 16.67 70.73 27.83 54.54 41.80 68.33 74.70 76.35 76.26 76.37
BianQue2-6B 7.00 13.87 15.13 39.67 14.76 21.35 10.33 13.93 4.73 3.30 3.85 3.25
DoctorGLM-6B 9.17 9.46 1.13 4.87 11.65 28.67 5.60 17.33 33.72 31.07 32.86 30.09
BenTsao-Llama-7B 0.00 0.11 0.80 2.47 4.50 9.79 3.33 9.33 23.00 21.12 22.69 21.46
ShenNong-TCM-LLM-7B 0.18 1.93 0.07 0.67 7.91 19.23 6.20 14.93 23.36 21.51 22.58 22.65
Taiyi-7B 22.57 22.26 19.87 33.20 23.71 31.52 33.47 54.60 43.05 44.80 43.93 45.16
@@
Zhongjing-13B 0.80 6.29 0.80 16.00 14.40 25.65 13.87 19.60 23.03 23.00 23.63 22.69
ChiMedGPT-13B 1.04 3.52 1.87 5.20 6.84 9.42 5.87 5.87 22.45 23.23 23.48 26.21
DISC-MedLLM-13B 2.55 3.54 0.13 1.80 9.86 16.53 14.07 12.13 37.09 35.57 37.53 35.44
@@
Table 3: Objective evaluation results of traditional Chinese medicine. DI denotes direct inference, CoT denotes chain-of-thought, ZS denotes zero-shot, and FS denotes few-shot.
4.2.2 Subjective Evaluation
@@
Table 4: Objective evaluation results of Chinese medical exam. Results marked with * are taken from the CMB benchmark leaderboard. ZS denotes zero-shot and FS denotes few-shot.

### v2

@@ 3.2 Supervised Fine-Tuning
TCM-Plan-SFT Medical Record Instruction 21M 13,409 70MB
@@
SFT is the core stage for applying the large language model of TCM to real-world scenarios, and the key phase in enabling the model’s conversational abilities. The quality and diversity of the instruction data are crucial to the SFT process [3]. To maximize the activation of the knowledge embedded in the pre-trained model and align it with downstream real-world TCM tasks, we constructed a high-quality and diverse instruction dataset, containing the following five types of data:
@@
TCM Pharmacopoeia Data. During the CPT stage, the raw text from the Pharmacopoeia of the People’s Republic of China (Part I: Traditional Chinese Medicine) was directly used to inject domain-specific knowledge into the model. To further enhance the model’s TCM instruction-following ability, we constructed two supervised instruction datasets—ChP-KnowledgeQA and ChP-PrescriptionWriting—based on this source. The ChP-KnowledgeQA dataset includes question-answer pairs covering topics such as the properties, identification, processing methods, taste, meridian tropism, functions, indications, dosage, and storage of TCM herbs. The ChP PrescriptionWriting dataset consists of QA pairs focused on composing and formulating TCM prescriptions. To ensure diversity, we manually designed a set of query templates when constructing these datasets. Together, these datasets constitute the ChP-TCM dataset, used exclusively in the supervised fine-tuning stage.
@@
Medical Record Data. Based on real hospital records, we constructed the TCMSD-DD-SFT, TCM-WM-Diff-SFT, and TCM-Plan-SFT instruction sets to improve the model’s performance on real-world tasks such as syndrome differentiation and disease diagnosis in TCM. The TCMSD-DD-SFT set is designed for TCM syndrome and disease differentiation, aiming to analyze a patient’s condition and syndrome type based on their chief complaint, admission details, medical history, and information gathered through TCM’s “Four Diagnoses” (inspection, listening and smelling, inquiry, and palpation). The TCM-WM-DiffDiag-SFT set includes differential diagnoses in both TCM and western medicine (WM) based on real hospital records. The TCM-Plan-SFT set focuses on providing detailed treatment plans based on a patient’s basic condition and diagnosis results. To construct these instruction sets, we manually created a variety of inquiry and response templates, and then employed a weighted random selection approach to generate instruction data from these templates, enhancing the quality and diversity of the instruction sets.
@@
Multi-turn Dialogue Data. To equip the model with conversational capabilities in the medical field, we utilized the DISC-Med-SFT dataset [17]. DISC-Med-SFT is a high-quality SFT dataset containing a large number of samples that were reconstructed from existing medical datasets. These data help the model learn medical knowledge, align its behavior with human preferences, and adapt to the distribution of real-world online medical dialogues.
@@
Exam Data. The MLEC-QA dataset is a large-scale Chinese biomedical multiple-choice question-answering dataset [38], sourced from the Chinese Medical Licensing Examination. It covers five subfields: TCM, integrative medicine (combining TCM and Western medicine), clinical medicine, public health, and basic clinical medicine. Based on this dataset, we constructed the MLEC-SFT instruction set to enhance the model’s comprehensive question-answering abilities in TCM and healthcare.
@@
General Domain Data. To prevent the loss of the model’s general capabilities, we mixed in some general domain data during the SFT process. MS-Agent-Bench [^4] https://www.modelscope.cn/datasets/iic/MSAgent-Bench is a general-purpose multi-domain SFT dataset, from which we randomly selected 65,000 samples. Self-Cognition is a custom instruction set we manually constructed to help the model develop self-awareness capabilities.
@@
The NLL loss for supervised fine-tuning based on dialogue or question answering can be defined as:
$\mathcal{L}_{\text{SFT}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}$ $\log P_{\theta}(y_{t}^{(i)}\mid x_{1}^{(i)},$ (2)
$\ldots,x_{T_{i}}^{(i)},y_{1}^{(i)},\ldots,y_{t-1}^{(i)}),$
where $x_{j}^{(i)}$ denotes input tokens (e.g., question or dialogue context) in the $i$-th example, whereas $y_{t}^{(i)}$ denotes the token at position $t$ in the response. And, $P_{\theta}(\cdot)$ denotes the probability of generating the token at position $t$ given the input context and previously generated response tokens, computed by the model with parameters $\theta$ [39, 40].
3.3 Model Training Setups
We train BianCang at two scales, with parameter sizes of 7 billion and 14 billion, based on Qwen2/2.5-7B and Qwen2.5-14B [36, 37], respectively. We employ full-parameter fine-tuning during both the continual pre-training and supervised fine-tuning stages. We use Scalable lightWeight Infrastructure for Fine-Tuning (SWIFT) [41] as the training framework.
To ensure training stability, we set the weight decay to 0.1 for the 14B model and 0.01 for the 7B model to prevent overfitting, the max gradient norm to 0.5 to prevent gradient explosion, and the warmup ratio to 0.05 for 14B model and 0.03 for 7B model to maintain smooth training. To balance training costs, we utilize parallel training, gradient checkpointing, fp16 precision, and a gradient accumulation strategy while limiting the model’s maximum response length to 4096 tokens. We reserve 5% of the training set for validation and use token-level accuracy to validate the model, saving the best checkpoint as the final model. Experiments are conducted on 8 A100 40G GPUs, with both continual pre-training and supervised fine-tuning performed over two epochs. The loss across all training stages successfully converges to a valid range. The continual pretraining process for the 7B model took approximately 174 hours, followed by 119 hours of supervised fine-tuning. In comparison, the 14B model required around 805 hours for continual pretraining and 605 hours for supervised fine-tuning.
4 Experiments
The evaluations are conducted from both objective and subjective perspectives. The objective evaluation includes 10 datasets, involving 3 tasks: TCM syndrome differentiation, TCM disease diagnosis, and exams. The subjective evaluation is conducted from 3 dimensions: professionalism, fluency, and safety, using the BC-Analytical dataset constructed by us.
@@ 4.1 Baselines
GPT-4o [5]: GPT-4o is a large-scale, multimodal model developed by OpenAI. It has demonstrated human-level performance across various professional and academic benchmarks. We conduct tests using the API service provided by OpenAI.
DeepSeek-V3 [3]: DeepSeek-V3 is a strong Mixture-of-Experts (MoE) language model with 671B total parameters and 37B activated for each token. It adopts Multi-head Latent Attention (MLA) and DeepSeekMoE architectures, achieving efficient inference and cost-effective training, and is pre-trained on 14.8 trillion diverse and high-quality tokens. DeepSeek-V3 outperforms other open-source models and achieves performance comparable to leading closed-source models.
DeepSeek-R1 [4]: DeepSeek-R1 is the first-generation reasoning model developed by DeepSeek AI, designed to enhance reasoning performance through multi-stage training and cold-start data. DeepSeek-R1 is suitable for tasks that require complex reasoning, such as mathematical problem solving, logical reasoning, and natural language understanding.
@@
BianQue2-6B [42]: BianQue-6B is a ChatGLM-based [43] LLM finetuned with the self-constructed health conversation dataset BianQueCorpus that consists of multiple turns of questioning and health suggestions polished by ChatGPT. BianQue2-6B expands upon the previous version by incorporating data from pharmaceutical instruction guides, medical encyclopedia knowledge, and ChatGPT distillation instructions, enhancing the model’s capabilities in providing recommendations and knowledge retrieval.
@@
DoctorGLM-6B [23]: A large-scale Chinese medical model based on ChatGLM-6B [43], fine-tuning on a large amount of medical instruction dataset.
@@
BenTsao-Llama-7B [15]: A Chinese medical LLM based on Chinese-LLaMA-7B [44], and fine-tuned on an 8K medical dialogue dataset.
@@
ShenNong-TCM-LLM-7B [31]: A large traditional Chinese medicine model based on Chinese-LLaMA-7B [44], fine-tuned on the self-constructed ChatMed-TCM dataset.
@@
Taiyi-7B [18]: A bilingual medical large language model designed for various biomedical tasks, trained on 140 Chinese and English biomedical datasets across 15 BioNLP tasks.
@@
WiNGPT2555https://github.com/winninghealth/WiNGPT2: A large language model developed by Weining Health for the medical domain, aimed at providing intelligent medical Q&A, diagnostic support, and medical knowledge services to improve the efficiency of diagnosis and the quality of healthcare services. WiNGPT2-7B-Base is based on Qwen-7B [45], while WiNGPT2-Llama3-8B-Base is based on LLaMA3-8B [46].
@@
HuatuoGPT2 [16]: HuatuoGPT2 employs an innovative domain adaptation method to significantly boost its medical knowledge and dialogue proficiency. We utilized two models, HuatuoGPT2-7B and HuatuoGPT2-13B, which are based on Baichuan2-7B-Base and Baichuan2-13B-Base [6], respectively.
@@
Sunsimiao-7B [47]: A Chinese medical large language model based on Qwen2-7B [36], fine-tuned on high-quality medical data. It has achieved excellent results in Chinese medical exams.
@@
QiZhen-CaMA-13B [^6] https://github.com/CMKRG/QiZhenGPT: A Chinese medical large language model based on CaMA-13B [48], and it was trained on the QiZhen medical knowledge base with a Chinese medical instruction set. It performs exceptionally well in Chinese medical scenarios.
@@
Zhongjing-13B [21]: A Chinese medical large language model that has undergone end-to-end training, from pre-training to reinforcement learning. In certain dialogue scenarios, it approaches the professional level of expert doctors. Zhongjing-13B is based on Ziya-LLaMA-13B-v1 [49], a general Chinese LLM with 13 billion parameters, trained based on LLaMA.
@@
ChiMedGPT-13B [25]: ChiMed-GPT is a Chinese medical large language model based on Ziya-LLaMA-13B-v2 [49]. It has undergone comprehensive pre-training, SFT, and RLHF.
@@
DISC-MedLLM-13B [17]: A large language model specifically designed for medical and healthcare conversational scenarios. DISC-MedLLM-13B is based on Baichuan-13B-Base [^7] https://github.com/baichuan-inc/Baichuan-13B and trained on the DISC-Med-SFT multi-turn dialogue dataset.
@@
Lingdan-13B [22]: A traditional Chinese medicine (TCM) large language model based on Baichuan2-13B-Base [6], trained on a rich dataset including TCM ancient books, textbooks, and the Chinese Pharmacopoeia.
@@
GLM4 [50]: GLM-4-9B is the open-source version of the latest generation of pre-trained models in the GLM-4 series launched by Zhipu AI.
@@
Qwen2/2.5 [37]: Qwen2 is a series of open-source multilingual pre-trained and instruction-tuned models based on the Transformer architecture, developed by the Qwen Team. These models have achieved outstanding results across various benchmark evaluations. Qwen2.5 represents a significant upgrade over Qwen2, optimized to meet specific requirements.
Table 3: Objective evaluation results of traditional Chinese medicine. DI denotes direct inference, CoT denotes chain-of-thought, ZS denotes zero-shot, and FS denotes few-shot. Notably, BianCang outperforms all baseline models with comparable scale (ranging from 6B to 14B), surpassing even GPT-4o and DeepSeek-V3 (671B) on all metrics. Moreover, compared to the reasoning model DeepSeek-R1 (671B), BianCang demonstrates significant advantages across multiple evaluations and achieves comparable results on most metrics.
@@
GPT-4o [5] 24.53 45.21 16.67 70.73 27.83 54.54 41.80 68.33 74.70 76.35 76.26 76.37
DeepSeek-V3 [3] 34.62 40.74 24.53 72.00 46.97 59.08 82.67 72.93 84.97 88.56 85.05 87.81
DeepSeek-R1 [4] 37.17 55.67 25.53 76.07 50.66 80.75 79.27 94.53 92.68 93.10 90.92 90.77
BianQue2-6B [42] 7.00 13.87 15.13 39.67 14.76 21.35 10.33 13.93 4.73 3.30 3.85 3.25
DoctorGLM-6B [23] 9.17 9.46 1.13 4.87 11.65 28.67 5.60 17.33 33.72 31.07 32.86 30.09
BenTsao-Llama-7B [15] 0.00 0.11 0.80 2.47 4.50 9.79 3.33 9.33 23.00 21.12 22.69 21.46
ShenNong-TCM-LLM-7B [31] 0.18 1.93 0.07 0.67 7.91 19.23 6.20 14.93 23.36 21.51 22.58 22.65
Taiyi-7B [18] 22.57 22.26 19.87 33.20 23.71 31.52 33.47 54.60 43.05 44.80 43.93 45.16
@@
Zhongjing-13B [21] 0.80 6.29 0.80 16.00 14.40 25.65 13.87 19.60 23.03 23.00 23.63 22.69
ChiMedGPT-13B [25] 1.04 3.52 1.87 5.20 6.84 9.42 5.87 5.87 22.45 23.23 23.48 26.21
DISC-MedLLM-13B [17] 2.55 3.54 0.13 1.80 9.86 16.53 14.07 12.13 37.09 35.57 37.53 35.44
@@
Table 4: Objective evaluation results of Chinese medical exam. Results marked with * are taken from the CMB benchmark leaderboard. ZS denotes zero-shot and FS denotes few-shot. BianCang demonstrates outstanding performance in modern medicine, surpassing all baseline models with comparable scale (ranging from 6B to 14B). Moreover, BianCang achieves results second only to DeepSeek-R1 (671B) in the comprehensive medical benchmark (CMB), highlighting its robust capabilities in modern medicine.
@@
4.2 Objective Evaluation Details and Datasets
For the TCM syndrome differentiation task, we used the TCMSD and TCMSD-BC datasets. The TCMSD dataset was adapted from the TCM-SD dataset [19], where we constructed queries and responses using fields from the test set of TCM-SD dataset. The TCMSD-BC dataset was built based on proprietary medical records. The TCM syndrome differentiation task is divided into two test modes: direct inference (DI) and chain of thought (CoT), with accuracy serving as the evaluation metric.
For the TCM disease diagnosis task, we used the TCMDD and TCMDD-BC datasets. Similarly, the TCMDD dataset was adapted from the test set of the TCM-SD dataset [19], while the TCMDD-BC dataset was constructed based on proprietary medical records. Like the syndrome differentiation task, the disease diagnosis task is evaluated using two test modes: DI and CoT, with accuracy as the evaluation metric. It should be noted that the inputs for TCM syndrome differentiation and disease diagnosis tasks tend to be lengthy. However, certain models (e.g., DoctorGLM and BianQue) are constrained by a limited maximum context length. To ensure fairness, both the DI and CoT evaluation modes for these tasks employ a zero-shot approach to avoid context overflow.
To assess the model’s mastery of TCM and medical knowledge, we applied the MLEC-QA [38] and Chinese Medical Benchmark (CMB [^8] https://cmedbenchmark.llmzoo.com/static/leaderboard.html) [32] datasets. The MLEC-QA dataset is divided into five subsets: traditional Chinese medicine, traditional Chinese medicine combined with western medicine, clinic, public health, and stomatology. The CMB dataset includes six categories: physician exam, nursing exam, pharmacist exam, medical technician exam, professional knowledge exam, and medical postgraduate entrance exam. The average score across these six categories is used as the final result. The examination tasks are evaluated in two test modes: zero-shot and few-shot, with accuracy as the evaluation metric.
We evaluated the model using the exact match (EM) method. Specifically, we first extracted the answer segment from the model’s response and then performed a word-by-word comparison between the extracted answer and the reference answer. A match was considered successful only if the two answers were entirely identical. Finally, we computed the proportion of successful matches to derive the accuracy, which served as the metric for evaluating the model’s performance.
4.3 Objective Evaluation Results
The objective evaluation results in the field of Traditional Chinese Medicine (TCM) are presented in Table 3. The experimental results demonstrate that BianCang achieves strong overall performance across syndrome differentiation, disease diagnosis, and TCM examination tasks. Most existing baselines exhibit weak analytical capabilities in syndrome differentiation and disease diagnosis. We identify three key factors contributing to this limitation: (1) their foundational models are often relatively weak, leading to insufficient general reasoning capacity; (2) TCM tasks require systematic knowledge and diagnostic experience that most models lack due to limited exposure to real-world TCM data; and (3) although some models utilize continual pre-training and instruction tuning, they often fail to effectively align these stages, resulting in inconsistencies between learned knowledge and fine-tuning objectives.
BianCang demonstrates robust analytical performance across both standard and custom datasets. For example, BianCang-Qwen2.5-7B-Instruct achieves an accuracy of 78.90% on the TCMSD dataset under direct inference (DI), outperforming the foundational model Qwen2.5-7B by over 48 percentage points, and surpassing the strongest TCM baseline Lingdan-13B by a similar margin. Under the chain-of-thought (CoT) setting, its accuracy rises to 82.10%, showing further improvements over both base and SOTA baselines. Comparable trends are observed on the TCMSD-BC, TCMDD, and TCMDD-BC datasets, confirming that BianCang generalizes well across different TCM task formats. Moreover, in the TCM examination tasks, BianCang achieves performance close to DeepSeek-R1 (670B), and notably achieves the highest accuracy in the Integrated Chinese and Western Medicine (MLEC-CWM) subtask, reflecting its comprehensive and structured grasp of both traditional and modern diagnostic paradigms.
It is worth noting that although BianCang achieves competitive results across the board, it is marginally outperformed by DeepSeek-R1 in three specific cases: TCMDD-BC (BianCang: 89.40%, DeepSeek-R1: 94.53%), MLEC-TCM (zero-shot) (92.39% vs. 92.68%), and MLEC-TCM (few-shot) (92.39% vs. 93.10%). These differences may be attributed to DeepSeek-R1’s massive parameter scale (670B), which allows it to better capture implicit reasoning patterns and general medical knowledge, particularly in few-shot and CoT-driven tasks. Nevertheless, BianCang attains highly competitive or superior performance with a much smaller 7B/14B model, underscoring the efficacy of our domain-specific training approach. Interestingly, we also observe that BianCang-Qwen2.5-7B-Instruct outperforms the larger 14B-Instruct variant in certain tasks, such as TCMSD (CoT: 82.10% vs. 75.54%) and TCMDD (CoT: 82.65% vs. 82.05%). This result may be due to larger models overinterpreting instructions, leading to unnecessary reasoning that reduces diagnostic accuracy. Smaller models tend to follow instructions more directly and focus on key information. Moreover, large models are more prone to training instability and forgetting during fine-tuning. Finally, we observe that the performance gap between DI and CoT inference in BianCang is considerably smaller than in other models—for instance, DeepSeek-R1 exhibits a 50.5-point gap on TCMSD-BC (DI: 25.53%, CoT: 76.07%), while BianCang-Qwen2.5-7B-Instruct exhibits a smaller gap of only 11 points (DI: 66.73%, CoT: 77.73%). This suggests that BianCang has internalized key diagnostic reasoning patterns through domain-specific training and does not rely heavily on external CoT prompting. In contrast, general-purpose models such as DeepSeek appear to depend more on explicit CoT reasoning to compensate for a lack of domain alignment, leading to significantly weaker performance in direct inference.
Additionally, we report more evaluation results on the Chinese medical examination in Table 4. As illustrated in Table 4, BianCang not only performs well in the TCM domain but also demonstrates outstanding performance in modern medicine, surpassing all baselines with comparable scale (ranging from 6B to 14B) in Clinic, Public Health, and Stomatology examinations. Moreover, BianCang achieves results second only to DeepSeek-R1 (671B) in the comprehensive medical benchmark (CMB), underscoring its robust capabilities across both TCM and modern medicine.